OpenStudy (anonymous):

Amelia is planning the trajectory of her next flight using the altitude & distance from Paris, France. She has determined her function to be f(x)=-4x+102. Based on the situation aboce, describe any necessary restrictions to the domain & range. Provide justification for the restrictions.

OpenStudy (anonymous):

That's my question! :D

OpenStudy (anonymous):

i saw yours but the person didn't really answer it

OpenStudy (anonymous):

Yeah :/ Thanks for helping!

OpenStudy (anonymous):

What did you put?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

as your answer

OpenStudy (anonymous):

did you answer the question?

OpenStudy (anonymous):

Oh, no I did not. I don't have a good answer!

OpenStudy (anonymous):

Oh..im screwed..lol

OpenStudy (anonymous):

Hahaha you need the answer too?

OpenStudy (anonymous):

yeah :s

OpenStudy (anonymous):

aww :P Yeah we're both screwed buddy!

OpenStudy (anonymous):

Haha..did you have it for algebra 2? FLVS

OpenStudy (anonymous):

Yep!!! Are you doing that?

OpenStudy (anonymous):

Yeah haha, did you get number 1 or 2?

OpenStudy (anonymous):

Well I got most of number one & number 2 yes lol Did you get number 5?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh what'd you get??

OpenStudy (anonymous):

It's the one with the graph right?

OpenStudy (anonymous):

I'll post it hold on, you should get me #1 ;)

OpenStudy (anonymous):

OKay!! :) Thanks!

OpenStudy (anonymous):

and i got 58.5 for #2 but i feel like its wrong

OpenStudy (anonymous):

For number 1, I. & II. are the same.

OpenStudy (anonymous):

I got 18.5

OpenStudy (anonymous):

heres #5: f(x): Slope=2, Y Intercept=5 g(x): Slope=-8, Y Intercept=-5 h(x): Slope=2, Y Intercept=-3 j(x): Slope=2, Y Intercept=-5 f(x), h(x), and j(x) are similar because they both have the same slope. g(x) and j(x) are similar because they both have the same Y Intercept. h(x) and f(x) are different because they do not share a similar Y Intercept with any other function.

OpenStudy (anonymous):

whatd you do for #2?

OpenStudy (anonymous):

and for #1 what were the steps?

OpenStudy (anonymous):

You're amazing! Thanks :) Here is #1 & #2: 1) 4x-14=2x+8 -2x -2x 2x-14=8 +14+14 2x=22 ----- 2 2 x=11 3x-21=12 +21+21 3x=33 ------ 3 3 x=11

OpenStudy (anonymous):

#2) f(x)=(3(24)+2)/4 f(x)=(72+2)/4 f(x)=18.5

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

now we gotta get that damn q

OpenStudy (anonymous):

damn question 4*

OpenStudy (anonymous):

Mhmm :) And yeah!! Only one!

OpenStudy (anonymous):

was your question 2 talking about Pedro creating a function to represent the number of assignments he has completed

OpenStudy (anonymous):

No, my number 2 was talking about a chick named Juliana texting

OpenStudy (anonymous):

how many texts she can send*

OpenStudy (anonymous):

Is your #5 talking about Jessie walking dogs?

OpenStudy (anonymous):

@bbbbbbbbbbbbbbbbbb ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

uh oh

OpenStudy (anonymous):

Noooooo :(

OpenStudy (anonymous):

my #5 was comparing and contrasting 4 functions

OpenStudy (anonymous):

Yeah here was my graph for #5 they gave me

OpenStudy (anonymous):

It said "Compare & contrast these four functions using complete sentences. Focus on the slope & y-intercept first, and then on any additional properties of each function."

OpenStudy (anonymous):

yeah thats the same as mine

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

hahaha oh phew! I got this answer btw :) I think... lol

OpenStudy (anonymous):

restrictions to the domain: X is less than or equal to her landing site. Restrictions to the range: 0 is less than Y is less than the ceiling of the aircraft.

OpenStudy (anonymous):

im still trying to figure it out..come on google!

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

ahh thanks!

OpenStudy (anonymous):

im gonna put that the domain and range can be all real number for my answer probably

OpenStudy (anonymous):

okay!! :)